OpenStudy (anonymous):

Find 3 solutions for each: 6y-21x=3 8y-5=28x 4y+9=x 5-2y=8x

OpenStudy (triciaal):

6y-21x =3 when x = 0 y = 3/6 = 1/2 when y = 0 x = -1/7 6y =3 + 21x y =7/2 x + 1/2